OpenStudy (anonymous):

how to prove this identity? tan A=cosec 2A-cot 2A

OpenStudy (anonymous):

=1/sin2A - cos2A/sin2A = (1-cos2A)/(sin2A) = 2 (sinA)^2/(2sinAcosA) = tanA